40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16399     Entry time: Wed Oct 13 15:36:38 2021
Author: Hang 
Type: Update 
Category: Calibration 
Subject: XARM OLTF 

We did a few quick XARM oltf measurements. We excited C1:LSC-ETMX_EXC with a broadband white noise upto 4 kHz. The timestamps for the measurements are: 1318199043 (start) - 1318199427 (end).

We will process the measurement to compute the cavity pole and analog filter poles & zeros later.

Attachment 1: Screenshot_2021-10-13_15-32-16.png  167 kB  Uploaded Wed Oct 13 16:37:37 2021  | Hide | Hide all
Screenshot_2021-10-13_15-32-16.png
ELOG V3.1.3-